# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

### v2

@@ Abstract
The rise of misinformation, exacerbated by Large Language Models (LLMs) like GPT and Gemini, demands robust fact-checking solutions, especially for low-resource languages like Vietnamese. Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency. We introduce SemViQA, a novel Vietnamese fact-checking framework integrating Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC). Our approach balances precision and speed, achieving state-of-the-art results with 78.97% strict accuracy on ISE-DSC01 and 80.82% on ViWikiFC, securing 1st place in the UIT Data Science Challenge. Additionally, SemViQA Faster improves inference speed 7× while maintaining competitive accuracy. SemViQA sets a new benchmark for Vietnamese fact verification, advancing the fight against misinformation. The source code is available at: https://github.com/DAVID-NGUYEN-S16/SemViQA.
@@ 1 Introduction
Although fact verification has been extensively studied in high-resource languages like English, with a primary focus on datasets such as FEVER Thorne et al. (2018), LIAR Wang (2017), PubHealthTab Akhtar et al. (2022), and TabFact Chen et al. (2020), extending these methods to low-resource languages like Vietnamese remains a significant challenge. Transformer-based approaches, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), have shown strong performance in fact verification, yet their adaptation to Vietnamese is still underexplored. For Vietnamese fact verification, ViNSV Tran et al. (2024) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval. However, SBERT’s 256-token input limitation leads to information loss in long-context claims, making it particularly ineffective for complex cases requiring multi-sentence reasoning or cross-referencing distant evidence. As a result, the system struggles with hard samples that demand deeper contextual understanding, ultimately reducing fact verification accuracy. In contrast, graph-based reasoning methods in English Zhong et al. (2020) have shown notable potential for semantic inference but are often hindered by high computational costs, restricting their real-world applicability. Moreover, traditional retrieval techniques like TF-IDF and BM25, though widely used, rely heavily on token frequency and exact keyword matching, making them less effective at capturing the nuanced semantics of complex claims. While some recent work employing large language models (LLMs) (Huo et al., 2023; Schimanski et al., 2024) has shown promise in addressing hard retrieval questions, these approaches typically require considerably more computing resources and longer processing times compared to conventional methods. Thus, a trade-off persists: models offering high accuracy tend to operate slowly, whereas faster models often compromise on accuracy.
To address the challenges in fact verification, we introduce SemViQA, a system designed to optimize both accuracy and efficiency. It comes in two variants: SemViQA Standard, which prioritizes accuracy, and SemViQA Faster, which significantly improves inference speed while maintaining strong performance. The system is built upon three core components:
@@
Data Processing: A specialized preprocessing pipeline designed to efficiently handle long-token sequences, a known limitation of Transformer-based models. Our approach preserves the semantic integrity of input claims while ensuring optimal tokenization for downstream processing.
@@
Semantic-based Evidence Retrieval (SER): A hybrid retrieval framework combining TF-IDF for efficient keyword-based matching and a Question Answering Token Classifier (QATC) for semantic-based evidence selection. This approach balances speed and accuracy by prioritizing TF-IDF for common cases while activating QATC selectively for hard samples, significantly improving precision while maintaining computational efficiency.
3.
Two-step Verdict Classification (TVC): A hierarchical classification strategy leveraging Focal Loss Lin et al. (2018) and Cross-Entropy Loss to enhance claim verification accuracy. This approach first determines whether a claim is Supported, Refuted, or Not Enough Information (NEI) and subsequently refines the decision boundary for hard-to-classify cases, reducing misclassification and improving robustness.
We evaluate SemViQA on two of the largest Vietnamese fact verification datasets, ISE-DSC01 and ViWikiFC Le et al. (2024). Our system achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC, outperforming existing baselines. These results highlight SemViQA’s potential to improve the accuracy of fact verification in Vietnamese, contributing to misinformation mitigation and enhancing transparency in digital content.
The remainder of this paper is structured as follows: Section 2 discusses related work. Section 3 details our proposed methodology, including SER, TVC, and QATC. Section 4 presents experimental results and performance evaluation. Finally, Section 5 concludes the paper and outlines future research directions.
@@ 2 Related Works
Advancements in Natural Language Processing (NLP) have significantly improved information verification and evidence extraction. Early methods, such as the Neural Semantic Matching Network (NSMN) Nie et al. (2018), leveraged BiLSTM Graves and Schmidhuber (2005) and WordNet-based features to enhance verification accuracy but struggled with complex sentence relations due to LSTM’s sequential nature.
The introduction of transformer models, notably BERT Devlin et al. (2019), transformed claim verification by enabling bidirectional contextual encoding Soleimani et al. (2020); Zhou et al. (2019); Malon (2018). While achieving state-of-the-art results on FEVER Aly et al. (2021); Lin et al. (2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), BERT’s input length limitation makes it less effective for long-context fact-checking.
Graph-based models have further advanced multi-step reasoning. Approaches such as GCN-based reasoning Zhong et al. (2020); Thorne et al. (2018) and AdMIRaL’s logic-driven document retrieval Aly and Vlachos (2022) improve evidence sufficiency but demand high computational resources.
For Vietnamese fact verification, research remains limited. ViNSV Tran et al. (2024) combined BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but struggled with complex reasoning due to static embeddings.
Building on prior work, SemViQA optimizes fact verification by integrating fast keyword-based retrieval (TF-IDF) with semantic reasoning (QATC) and a hierarchical classification strategy, ensuring high accuracy, fast inference, and scalability for real-world applications.
@@ 3 SemViQA - Semantic Vietnamese Question Answering
Our proposed method, SemViQA, employs a three-stage pipeline to enhance fact verification, as illustrated in Figure 1. The process begins with a preprocessing stage, where input data is restructured to ensure compatibility with retrieval and classification models. Subsequently, evidence retrieval is performed using TF-IDF for straightforward cases, while the Question Answering Token Classifier (QATC) refines evidence selection for more complex claims. Finally, in the claim classification stage, a hierarchical approach integrates Cross-Entropy Loss and Focal Loss Lin et al. (2018) to improve both accuracy and robustness.
3.1 Data Processing
Before implementing data processing solutions, we conducted a thorough analysis of the input data to understand its characteristics and challenges. One of the prominent issues we encountered is the token limit of current Vietnamese BERT models, as clearly illustrated in Figure 2.
During data processing, we observed that the input context has a considerable length, with the longest context extending up to 4803 tokens. This presents a challenge for data processing given the limitations of 256 tokens for classification and 512 tokens for the Question Answering model.
[…]
From a long context, we proceed to split it into smaller segments. We then slide through each segment and check the number of tokens in each segment. If the token count does not exceed 400, we append the segment to the subcontext. When the subcontext has a token count greater than 400 (see Figure 3), we check if the evidence sentence is contained within the subcontext. If it is, then the subcontext is considered to contain that evidence. Otherwise, if it is not present, the evidence for that subcontext is assigned as an empty string. This process allows us to create new data samples with contexts containing fewer than 512 tokens, which is beneficial for training the model more effectively. This ensures that no segment is omitted, unlike stride-based cutting methods, which may result in evidence sentences being split between segments within a complete paragraph.
@@ 3.2 Semantic-based Evidence Retrieval (SER)
In addressing this problem, focusing on evidence retrieval is a crucial and decisive part. Accurate classification can only be performed when the evidence is correct. After analyzing and training on the data, we found that methods like TF-IDF and BM25 are effective for simpler samples but struggle with more complex samples.
Therefore, we have divided the evidence retrieval process into two parts:
Part 1: Using TF-IDF: We begin by segmenting the context into smaller passages and pairing each segment with the claim to generate structured data samples, which then undergo a preprocessing pipeline including noise removal and tokenization using ViTokenizer 222https://github.com/trungtv/pyvi to standardize the input. A key observation is that TF-IDF performs well on simple claims, particularly those classified as refuted, but struggles with claims requiring deeper reasoning, as it relies solely on keyword matching without capturing semantic relationships. To address this limitation, we introduce an enhanced strategy to enrich contextual information, where segments containing fewer than 60% of the claim’s token count are merged with the preceding segment to supplement missing information. This is particularly beneficial for short or incomplete sentences where crucial words may be absent from the claim, leading to improved evidence retrieval accuracy. After retrieving potential evidence using TF-IDF, we rank the retrieved segments and apply a Confidence Threshold to classify them into easy and challenging cases, as detailed in Appendix C. While simple cases can be resolved directly using TF-IDF, more complex cases are forwarded to Part 2 for further analysis, ensuring higher classification accuracy.
Part 2: Using the Question Answering Token Classifier Model: The subcontext is also used for model QA inference instead of inferring the full context. In this process, three situations can occur: The first, when multiple subcontexts generate different responses, we interpret this as the model exhibiting confusion in handling evidence responses. In this case, we collect all retrieved evidence and apply TF-IDF to select the most appropriate one. The second, when only one piece of evidence appears across multiple subcontexts, we interpret this as the model being highly confident in its prediction. The third, when the model does not retrieve any evidence, we use TF-IDF to extract relevant evidence as input for our TVC method. By combining these methods, we have achieved the best results in accurately and reliably retrieving evidence.
@@ 3.3 Two-step Verdict Classification (TVC)
Our input data consists of claim-evidence pairs, where the claim is the statement that needs to be verified, and the evidence comprises information used to support (supported), refute (refuted), or indicate insufficient information (NEI) regarding the claim. To address this problem, we undertake two main tasks as follows:
3.4 Three-Class Classification
The initial step involves classifying each claim-evidence pair into one of three categories: Supported, Refuted, or Not Enough Information (NEI). We define a three-class classification function as follows:
where denotes the predicted probability distribution over the three classes, and represents a neural network model based on an advanced BERT-derived architecture, fine-tuned for this task. The predicted label is determined by:
where is the class with the highest predicted probability. The model is optimized using Cross-Entropy Loss, ensuring that aligns closely with the ground-truth distribution.
3.5 Binary Classification
When the three-class model does not predict NEI (i.e., ), a secondary binary classification step refines the decision between Supported and Refuted. We define a binary classification function:
where represents the predicted probability distribution over the two classes, and is a neural network based on a fine-tuned BERT-derived architecture, specifically trained for binary classification. The predicted label is given by:
where corresponds to the class with the highest probability. This model is optimized using Focal Loss Lin et al. (2018) to address class imbalance and enhance performance on difficult examples.
3.6 Final Decision Rule
The final predicted class, denoted , is derived by integrating the outputs of the three-class and binary classification models. When the three-class model predicts , the decision is immediate. In all other cases, a comparison is made between the maximum probability from the three-class model (excluding the NEI class) and that from the binary model. This process is formalized by the following decision rule:
where represents the highest probability assigned by the three-class model to either the Supported or Refuted class, and denotes the corresponding maximum probability from the binary model.
This approach leverages the complementary strengths of both models: the three-class model provides a broad categorization, including the critical NEI class, while the binary model offers refined discrimination between Supported and Refuted claims. By prioritizing the three-class prediction when its confidence exceeds that of the binary model, the method ensures robust handling of complex claims while maintaining high accuracy for well-defined cases.
3.7 Question Answering Token Classifier (QATC)
Typical QA models are trained only to predict the start and end positions of an answer. Here, we extend this by introducing a Token Classification task, where we assign a label of 1 to tokens within the answer and 0 otherwise. This enhances the model’s ability to identify relevant tokens for answer extraction.
Procedure: We input the context and claim into the model, extract token features, and apply two linear layers to predict answer start and end probabilities. The default QA loss is the Cross Entropy loss:
where is the number of classes, is the true label, and is the softmax probability for the class.
3.7.1 Token Classification Task
Inspired by Rationale Tagging Ju et al. (2019), we predict whether each token is part of the answer (label 1) or not (label 0). For NEI (Not Enough Information) cases, all tokens are labeled 0. To achieve this, we introduce an additional fully connected (FC) layer that computes classification probabilities:
where is the output representation of token from RoBERTa Zhuang et al. (2021); Aly et al. (2021). The loss for this task is the Binary Cross Entropy loss:
3.7.2 Rationale Regularization Loss
To encourage sparsity and continuity in rationale predictions, we introduce a Rationale Regularization Loss , which consists of:
1. Sparsity Loss: Encourages the model to minimize the number of tokens selected as rationales:
where is a mask ensuring that the loss is computed only on non-padding tokens.
2. Continuity Loss: Encourages smoother rationale selection by minimizing sudden jumps between selected tokens:
The total rationale regularization loss is:
3.7.3 Combined Loss Function
The final objective function integrates all components:
where controls the weight of the rationale tagging loss, and , are hyperparameters for regularization.
@@ 4.1 Dataset
ISE-DSC01, used in the UIT Challenge 2023, is designed for AI evaluation in information extraction and verification. It comprises training, public test, and private test sets with diverse claim-context pairs, posing challenges in evidence retrieval and fact-checking.
ViWikiFC Le et al. (2024) was developed to combat misinformation, featuring over 20,000 claims from Wikipedia. Unlike ISE-DSC01, it explicitly includes evidence for the “nei” (Not Enough Information) label, enhancing real-world applicability.
| ISE-DSC01 | ViWikiFC | |
| Train | 37,967 | 16,738 |
| Dev | 4,794 | 2,090 |
| Test | 5,396 | 2,091 |
Table 1 summarizes the datasets. For ViWikiFC, experiments were conducted on all three sets: training, public test, and private test.
| Method | Parameter | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | ||
| TF-IDF | InfoXLMlarge | 560M | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
| XLM-Rlarge | 560M | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
| Ernie-Mlarge | 560M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
| BM25 | InfoXLMlarge | 560M | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
| XLM-Rlarge | 560M | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
| Ernie-Mlarge | 560M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
| SBert | InfoXLMlarge | 838M | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
| XLM-Rlarge | 838M | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
| Ernie-Mlarge | 838M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
| QA-based approaches | VC | ||||||||||
| ViMRClarge | InfoXLMlarge | 1120M | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
| XLM-Rlarge | 1120M | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
| Ernie-Mlarge | 1120M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
| InfoXLMlarge | InfoXLMlarge | 1120M | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
| XLM-Rlarge | 1120M | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
| Ernie-Mlarge | 1120M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
| LLM | |||||||||||
| Qwen2.5-1.5B-Instruct | 1.5B | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
| Qwen2.5-3B-Instruct | 3B | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
| LLM | VC | ||||||||||
| Qwen2.5-1.5B-Instruct | InfoXLMlarge | 2B | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
| XLM-Rlarge | 2B | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
| Ernie-Mlarge | 2B | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
| Qwen2.5-3B-Instruct | InfoXLMlarge | 3.5B | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
| XLM-Rlarge | 3.5B | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
| Ernie-Mlarge | 3.5B | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
| SER Faster (ours) | TVC (ours) | ||||||||||
| TF-IDF + ViMRClarge | Ernie-Mlarge | 1680M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
| TF-IDF + InfoXLMlarge | 1680M | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
| SER (ours) | TVC (ours) | ||||||||||
| TF-IDF + ViMRClarge | InfoXLMlarge | 1680M | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
| XLM-Rlarge | 1680M | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
| Ernie-Mlarge | 1680M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
| TF-IDF + InfoXLMlarge | InfoXLMlarge | 1680M | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
| XLM-Rlarge | 1680M | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
| Ernie-Mlarge | 1680M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |

### v3

@@ Abstract
Recent advances in LLMs have accelerated both information generation and misinformation, especially in low-resource languages like Vietnamese, motivating robust fact-checking systems. Existing methods struggle with semantic ambiguity, homonyms, and complex linguistic structures, often trading accuracy for efficiency. We introduce SemViQA, a novel Vietnamese fact-checking framework integrating Semantic-based Evidence Retrieval (SER) and Two-step Verdict Classification (TVC). Our approach balances precision and speed, achieving state-of-the-art results with 78.97% strict accuracy on ISE-DSC01 and 80.82% on ViWikiFC, securing 1st place in the UIT Data Science Challenge. Additionally, SemViQA Faster improves inference speed 7× while maintaining competitive accuracy. SemViQA sets a new benchmark for Vietnamese fact verification, advancing the fight against misinformation. The source code is available at: https://github.com/DAVID-NGUYEN-S16/SemViQA.
@@ 1 Introduction
Although fact verification has been widely studied in high-resource languages like English, applying these methods to low-resource languages such as Vietnamese remains a significant challenge. Transformer-based models, including BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), have demonstrated strong performance but their adaptation to Vietnamese is still limited. ViNSV Tran et al. (2024b) employs BM25 and SBERT Reimers and Gurevych (2019a) for evidence retrieval but suffers from SBERT’s 256-token input constraint, making it ineffective for complex, long-context claims. Graph-based reasoning methods Zhong et al. (2020) offer promising semantic inference but are often computationally expensive. Traditional retrieval methods like TF-IDF and BM25, while efficient, rely heavily on exact keyword matching, limiting their ability to capture nuanced semantics. Recent large language model (LLM) approaches Huo et al. (2023); Schimanski et al. (2024) show potential but typically require substantial computational resources, creating a trade-off between speed and accuracy.
To address these challenges, we propose SemViQA, a Vietnamese fact-checking framework that balances semantic accuracy and computational efficiency. As shown in Figure 1, SemViQA comprises three key components:
@@
Semantic-based Evidence Retrieval (SER): Includes a preprocessing step that efficiently handles long-token contexts by splitting them into manageable subcontexts (e.g., 400 tokens). It combines fast TF-IDF retrieval with selective Question Answering Token Classification (QATC) to strike a balance between speed and semantic accuracy.
@@
Two-step Verdict Classification (TVC): Employs a hierarchical classification strategy with both three-class and binary classification stages to enhance robustness and improve performance on challenging claim verification tasks.
SemViQA achieves 78.97% strict accuracy on ISE-DSC01 111https://codalab.lisn.upsaclay.fr/competitions/15497#results and 80.82% on ViWikiFC Le et al. (2024), outperforming existing baselines (see Table 1). These results validate SemViQA’s potential to enhance Vietnamese fact verification, supporting misinformation mitigation and improved transparency.
The rest of this paper is organized as follows: Section 2 reviews related work, Section 3 presents the methodology, Section 4 reports experimental results, and Section 5 concludes the paper with future directions.
@@ 2 Related Works
Advances in Natural Language Processing (NLP) have driven rapid progress in fact verification and evidence extraction. Early BiLSTM‑based models such as the Neural Semantic Matching Network (NSMN) Nie et al. (2018)—augmented with WordNet features improved accuracy but struggled with complex sentence relations due to sequential limitations Graves and Schmidhuber (2005). Transformer models, notably BERT Devlin et al. (2019), introduced bidirectional contextual encoding and achieved state‑of‑the‑art results on FEVER Soleimani et al. (2020); Zhou et al. (2019); Malon (2018); Aly et al. (2021); Lin et al. (2024); Yuan and Vlachos (2024); DeHaven and Scott (2023), yet their input‑length cap hinders long‑document fact‑checking. Graph‑based reasoning further enhances multi‑hop verification GCN variants Zhong et al. (2020); Thorne et al. (2018) and AdMIRaL’s logic‑driven retrieval Aly and Vlachos (2022) boost evidence sufficiency but at considerable computational cost. Vietnamese research is sparse: ViNSV Tran et al. (2024b) pairs BM25 with SBERT Reimers and Gurevych (2019a) yet falters on complex reasoning due to static embeddings.
TF-IDF remains the industry standard for document retrieval due to its speed, simplicity, interpretability, and ability to handle long contexts effectively, making it suitable for rapid and scalable retrieval Reddy et al. (2018); Qaiser and Ali (2018); Li (2021); Azevedo et al. (2022). However, its reliance on surface-level keyword matching limits its capability to handle paraphrases, contextual nuances, and multi-hop reasoning, reducing recall accuracy on complex queries. Ensemble learning Hannichenko et al. (2023); Wang et al. (2021); Liu et al. (2024); Ganaie et al. (2022) mitigates individual model weaknesses by aggregating diverse architectures and training signals, yielding robust gains. Building on these insights, SemViQA fuses fast TF‑IDF retrieval, semantic reasoning via QATC, and hierarchical classification, delivering high accuracy, low latency, and practical scalability for Vietnamese fact verification.
@@ 3 SemViQA - Semantic Vietnamese Question Answering
We formulate Vietnamese fact verification as a multi-output classification task, where the input is a pair , with being a claim and its corresponding context paragraph or document. The objective is to (i) identify the most relevant evidence sentence(s) from and (ii) predict the veracity label of the claim as one of three categories: Supported, Refuted, or Not Enough Information (NEI). To address challenges such as long input sequences and semantic ambiguity, we propose SemViQA, a three-stage framework consisting of data pre-processing, semantic-based evidence retrieval, and two-step verdict classification. An overview of the architecture is shown in Figure 2, with detailed descriptions provided in the following subsections.
3.1 Evidence Extraction via Question Answering with Token Classification
3.1.1 Data Processing
To effectively support evidence retrieval and claim classification, we apply distinct preprocessing strategies to the context based on the specific requirements of each downstream task (see Appendix A). A key challenge arises from the considerable length of many context passages, frequently exceeding the token limits of Vietnamese BERT-based models. Figure 3 illustrates our approach for handling long input contexts. First, the context is segmented into individual sentences. Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens. Each completed subcontext is then processed by the QATC model to identify potential evidence. The next subcontext begins from the subsequent sentence, and this process continues until all sentences are processed. However, processing subcontexts sequentially can be time-consuming. Therefore, we developed SemViQA Faster, which batches and processes subcontexts in parallel, significantly accelerating the retrieval process.
3.1.2 Question Answering with Token Classification (QATC)
Traditional Question Answering models typically predict the start and end positions of an answer span. In our framework, we enhance this approach by introducing a token-level classification objective, enabling the model to focus explicitly on tokens within evidence sentences in the context. This dual formulation provides improved supervision for evidence extraction. Drawing inspiration from rationale tagging Ju et al. (2019), we treat token labeling as a binary classification task: tokens within evidence sentences receive a label of 1, and all other tokens receive a label of 0. In cases marked NEI (Not Enough Information), every token is labeled as 0. We employ a feed-forward classification layer on the token representations:
where is the contextual representation of token , are learnable weights in the neural network and denotes the sigmoid function. The loss for this task is the Binary Cross-Entropy (BCE) loss:
@@ 3.2 Semantic-based Evidence Retrieval (SER)
Accurate claim verification requires reliable evidence. To improve both efficiency and robustness, we adopt a two-stage evidence retrieval strategy combining TF-IDF with a QATC.
Stage 1: TF-IDF-based Retrieval. We segment the context into smaller passages and pair each with the claim . Preprocessing includes noise removal and tokenization using ViTokenizer222https://github.com/trungtv/pyvi. TF-IDF is effective for simple claims particularly refuted ones but struggles with semantically complex cases due to its reliance on keyword overlap. To enrich short segments (i.e., those with fewer than 60% of ’s tokens), we merge them with preceding segments to improve evidence completeness. Retrieved segments are then ranked, and a confidence threshold is applied to identify easy cases (handled by TF-IDF) and hard cases (passed to QATC).
Stage 2: QATC-based Refinement. For complex cases, QATC is applied to segmented subcontexts rather than the full context due to the input length limitation of BERT models. The detailed processing approach is described in Section 3.1.1. At this time, we consider three scenarios: (1) If multiple subcontexts yield conflicting answers, we collect all predicted spans and re-rank them using TF-IDF. (2) If a single evidence span consistently appears, it is selected directly. (3) If no evidence is found, fallback to TF-IDF is used.
This hybrid approach balances speed and semantic accuracy, improving evidence selection for downstream verdict classification. Examples are provided in Appendix E.
@@ 3.3 Two-step Verdict Classification (TVC)
We adopt a two-stage classification framework to enhance claim verification robustness and mitigate label imbalance, especially the overrepresentation of NEI.
Stage 1: Three-Class Classification.
Given a claim-evidence pair , a BERT-based model predicts a probability distribution over three labels: Supported, Refuted, and Not Enough Information (NEI):
This step is optimized using Cross-Entropy Loss.
Stage 2: Binary Classification.
If , we apply a refined binary classifier to distinguish between Supported and Refuted:
This model uses Focal Loss Lin et al. (2018) to address class imbalance.
Final Prediction Rule.
The final verdict where is determined by comparing the confidence scores from both classifiers. Here, represents the index of the predicted label, where each index corresponds to a specific class description.
where , represents the highest probability.
This hybrid strategy allows the three-class model to handle general cases, especially detecting NEI early, while the binary model specializes in distinguishing difficult SUP/REF cases.
3.4 SemViQA Pipeline System
We now describe the full SemViQA pipeline, as illustrated in Figure 2. First, we prepare input for TF-IDF by splitting the context paragraph into sentences, then concatenating each sentence with the claim . We calculate the matching score for each sentence and select the one with the highest probability. If this score exceeds the threshold , we directly use this sentence as the evidence. If the score is below , we proceed to prepare input for the QATC model. We segment into subcontexts (as detailed in Section 3.1.1). Each subcontext is sequentially processed by the QATC model. If QATC identifies zero or multiple candidate evidence spans, we collect all predicted spans and re-rank them using TF-IDF. If QATC finds exactly one candidate evidence span, we confidently select it as the final evidence. Finally, we move to the two-step verdict classification (TVC) stage. We prepare input for TVC by concatenating the claim with the final evidence. We first apply the three-class model. If it predicts NEI, the process ends. Otherwise, we use an ensemble method to combine the weights from both the three-class and binary models to make the final prediction. All coefficients used in our system are provided in the ablation study, as shown in Appendix C.
@@ 4.1 Dataset
We use two Vietnamese fact verification datasets: ISE-DSC01 from the UIT Challenge 2023 and ViWikiFC Le et al. (2024), which contains over 20,000 Wikipedia-based claims, including annotated evidence for the “nei” label. Dataset details are provided in Appendix A.1.
| Method | ViWikiFC | ISE-DSC01 | Avg Strict Acc | |||||||
| ER | VC | Strict Acc | VC Acc | ER Acc | Time (s) | Strict Acc | VC Acc | ER Acc | Time (s) | |
| Traditional Baselines | ||||||||||
| TF-IDF | InfoXLM | 75.56 | 82.21 | 90.15 | 131 | 73.59 | 78.08 | 76.61 | 378 | 74.58 |
| XLM-R | 76.47 | 82.78 | 90.15 | 134 | 75.61 | 80.50 | 78.58 | 366 | 76.04 | |
| Ernie-M | 75.56 | 81.83 | 90.15 | 144 | 78.19 | 81.69 | 80.65 | 403 | 76.88 | |
| BM25 | InfoXLM | 70.44 | 79.01 | 83.50 | 130 | 72.09 | 77.37 | 75.04 | 320 | 71.27 |
| XLM-R | 70.97 | 78.91 | 83.50 | 132 | 73.94 | 79.37 | 76.95 | 333 | 72.46 | |
| Ernie-M | 70.21 | 78.29 | 83.50 | 141 | 76.58 | 80.76 | 79.02 | 381 | 73.40 | |
| SBert | InfoXLM | 74.99 | 81.59 | 89.72 | 195 | 71.20 | 76.59 | 74.15 | 915 | 73.10 |
| XLM-R | 75.80 | 82.35 | 89.72 | 194 | 72.85 | 78.78 | 75.89 | 835 | 74.33 | |
| Ernie-M | 75.13 | 81.44 | 89.72 | 203 | 75.46 | 79.89 | 77.91 | 920 | 75.30 | |
| QA-based Approaches | ||||||||||
| ViMRC | InfoXLM | 77.28 | 81.97 | 92.49 | 3778 | 54.36 | 64.14 | 56.84 | 9798 | 65.82 |
| XLM-R | 78.29 | 82.83 | 92.49 | 3824 | 53.98 | 66.70 | 57.77 | 9809 | 66.14 | |
| Ernie-M | 77.38 | 81.92 | 92.49 | 3785 | 56.62 | 62.19 | 58.91 | 9833 | 67.00 | |
| InfoXLM | InfoXLM | 78.14 | 82.07 | 93.45 | 4092 | 53.50 | 63.83 | 56.17 | 10057 | 65.82 |
| XLM-R | 79.20 | 83.07 | 93.45 | 4096 | 53.32 | 66.70 | 57.25 | 10066 | 66.26 | |
| Ernie-M | 78.24 | 82.21 | 93.45 | 4102 | 56.34 | 62.36 | 58.69 | 10078 | 67.29 | |
| LLMs | ||||||||||
| Qwen2.5-1.5-Instruct | 51.03 | 65.18 | 78.96 | 7665 | 59.23 | 66.68 | 65.51 | 19780 | 55.13 | |
| Qwen2.5-3B-Instruct | 44.38 | 62.31 | 71.35 | 12123 | 60.87 | 66.92 | 66.10 | 31284 | 52.63 | |
| Qwen2.5-1.5-Instruct | InfoXLM | 66.14 | 76.47 | 78.96 | 7788 | 64.40 | 68.37 | 66.49 | 19970 | 65.27 |
| XLM-R | 67.67 | 78.10 | 78.96 | 7789 | 64.66 | 69.63 | 66.72 | 19976 | 66.17 | |
| Ernie-M | 66.52 | 76.52 | 78.96 | 7794 | 65.70 | 68.37 | 67.33 | 20003 | 66.11 | |
| Qwen2.5-3B-Instruct | InfoXLM | 59.88 | 72.50 | 71.35 | 12246 | 65.72 | 69.66 | 67.51 | 31477 | 62.80 |
| XLM-R | 60.74 | 73.08 | 71.35 | 12246 | 66.12 | 70.44 | 67.83 | 31483 | 63.43 | |
| Ernie-M | 60.02 | 72.21 | 71.35 | 12251 | 67.48 | 70.77 | 68.75 | 31512 | 63.80 | |
| Ours: SER Faster + TVC | ||||||||||
| ViMRC | Ernie-M | 79.44 | 82.93 | 94.60 | 410 | 78.32 | 81.91 | 80.26 | 995 | 78.88 |
| InfoXLM | 79.77 | 83.07 | 95.03 | 487 | 78.37 | 81.91 | 80.32 | 925 | 79.07 | |
| Ours: Full SER + TVC | ||||||||||
| ViMRC | InfoXLM | 80.25 | 83.84 | 94.69 | 2731 | 75.13 | 79.54 | 76.87 | 5191 | 77.69 |
| XLM-R | 80.34 | 83.64 | 94.69 | 2733 | 76.71 | 81.65 | 78.91 | 5219 | 78.53 | |
| Ernie-M | 79.53 | 82.97 | 94.69 | 2733 | 78.97 | 82.54 | 80.91 | 5225 | 79.25 | |
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 3860 | 75.13 | 79.60 | 76.87 | 5175 | 77.91 |
| XLM-R | 80.82 | 83.88 | 95.31 | 3843 | 76.74 | 81.71 | 78.95 | 5200 | 78.78 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 3891 | 78.97 | 82.49 | 80.91 | 5297 | 79.52 | |